OpenStudy (anonymous):

find all imaginary and real zeros g(x)=x^3+9x^2+25x+25 explain please

myininaya (myininaya):

Did you check for any rational zeros yet?

myininaya (myininaya):

What are factors of 25: \[\pm1,\pm5,\pm25\] See if any of these numbers I listed are zeros :)

OpenStudy (anonymous):

-5

OpenStudy (anonymous):

only zero is -5

myininaya (myininaya):

\[g(-5)=(-5)^3+9(-5)^2+25(-5)+25=0\] So that means x=-5 is a zero So x+5 is a factor (since x=-5 means we had x+5=0 ) But anyways if x=-5 is a solution then you can find the other zeros by synthetic division + quadratic formula

myininaya (myininaya):